OpenStudy (anonymous):

solve using two equations with two variables Sarah took the advertising department from her company on a round trip to Chicago to meet with a potential client. Including Sarah, a total of 14 people took the trip. She was able to purchase coach tickets for $250 and first-class tickets for $1150. She used her total budget for airfare for the trip, which was $10,700. How many first-class tickets did she buy? How many coach tickets did she buy?

OpenStudy (anonymous):

I need help in setting up my equations.

OpenStudy (anonymous):

I haven't done word problems like this, but this would be how I envision the setup: You have x amount of people flying coach, and y amount of people flying first-class, with the total cost being $10,700. 250x+1150y=10700. Now, because x and y have to add up to be 14, x+y=14. Now you can either use substitution; matrix manipulation; elimination, what have you, to solve this.

OpenStudy (anonymous):

I set it up like that and I am stuck solving the equation. I am coming up with a bunch of fractions.

OpenStudy (anonymous):

Okay, try using substitution. In the second equation, set x = 14-y. Plug (14-y) into your first equation, in place of x, and you'll get a clean value for y. Then, you know how many first-class; which leads you to how many coach tickets.

OpenStudy (anonymous):

That's it. Thanks!!

OpenStudy (anonymous):

Glad to help. :)